OpenStudy (josedavid):

@amistre64 @Alchemista @AuroraB @AngelPSP @AndrewTheCookie @ashhhrodriguez @satellite73 @saifoo.khan @sammixboo @sleepyjess @Skyz @sweetburger @dan815 @Demonx341 @dmndlife24 @Donblue22 @damara1213 @danish071996 @fallenangelorchid @fantasiablueskys @faye0422 @fiat21 @flydiva @Gravity_Dreams @Ganpat @gazed @GinnyParis13 @G-6a6y

OpenStudy (sleepyjess):

Don't mass tag please

OpenStudy (fallenangelorchid):

And your question is?

OpenStudy (anonymous):

@josedavid

OpenStudy (anonymous):

What's your question?

OpenStudy (josedavid):

sorry gust about that is that my comp froze

OpenStudy (josedavid):

guys

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (fallenangelorchid):

I'm not going to give you the answer but... a1 is first term, d is common difference

OpenStudy (fallenangelorchid):

Thats just a hint.

OpenStudy (josedavid):

okay thanks

OpenStudy (anonymous):

@suckerofmath

OpenStudy (anonymous):

Okay I think to be able to solve this problem you have to find the common difference, i.e. what number you would add to the previous number to achieve the next number in the sequence.

OpenStudy (michele_laino):

your sequence is a arithmetic progression, with first term=4 and constant=3, now the n-term is: \[a _{n}=a _{1}+(n-1)*c=4+(n-1)*3=4+3n-3=3n+1=f(n)\]